# Preference-based Centrality and Ranking in General Metric Spaces

### v1

@@ Abstract
Assessing centrality or ranking observations in multivariate or non-Euclidean spaces is challenging because such data lack an intrinsic order and many classical depth notions lose resolution in high-dimensional or structured settings. We propose a preference-based framework that defines centrality through population pairwise proximity comparisons: a point is central if a typical draw from the underlying distribution tends to lie closer to it than to another. This perspective yields a well-defined statistical functional that generalizes data depth to arbitrary metric spaces. To obtain a coherent one-dimensional representation, we study a Bradley–Terry–Luce projection of the induced preferences and develop two finite-sample estimators based on convex M-estimation and spectral aggregation. The resulting procedures are consistent, scalable, and applicable to high-dimensional and non-Euclidean data, and across a range of examples they exhibit stable ranking behavior and improved resolution relative to classical depth-based methods.
Keywords: multivariate ranking; statistical depth; pairwise comparison; Bradley–Terry–Luce model; general metric spaces
@@ 1 Introduction
Centrality is a fundamental concept in statistics and the applied sciences. It underlies representative point selection, outlier detection, risk stratification, and the construction of interpretable orderings. In univariate settings, the real line provides a canonical order, so centrality and ranking are naturally defined through ranks, quantiles, and order statistics. In contrast, modern data are frequently multivariate or live on non-Euclidean spaces such as manifolds, networks, graphs, and learned embedding spaces, where no intrinsic notion of order is available. In these settings, defining population-level centrality and estimating a stable, interpretable ordering from finite samples remain challenging problems.
A broad literature has sought to generalize ranking and centrality beyond the univariate case. Classical approaches include componentwise and spatial ranks (Hallin and Puri, 1995; Puri and Sen, 1966; Chaudhuri, 1996; Oja, 2010), depth-induced orderings (Liu and Singh, 1993; Zuo and Serfling, 2000), Mahalanobis and metric-based constructions (Hallin and Paindaveine, 2002, 2006; Pan et al., 2018), and more recent transport- or graph-based formulations (Deb and Sen, 2023; Zhou and Chen, 2023). Statistical depth, in particular, has emerged as a principled framework for multivariate centrality, formalizing center–outward ordering through axioms such as affine equivariance, maximality, monotonicity, and robustness (Tukey, 1974; Liu, 1990; Vardi and Zhang, 2000; Zuo, 2003; Parzen, 1962). Despite their success, many existing depth notions rely on linear or convex structure and can lose discriminative power or become computationally burdensome in high dimensions or in complex geometries, where distance concentration, hubness, and anisotropy obscure purely geometric notions of centrality (Radovanovic et al., 2010; Feldbauer and Flexer, 2019). More broadly, when only a general dissimilarity is available, it is often unclear what population object a given surrogate score targets, and how pairwise information across the sample should be coherently aggregated into a global ordering.
@@
Intuitively, $r_{F}(y)$ is the probability that $y$ “wins” against a random opponent when judged by a random reference point. This functional provides a natural, population-level notion of centrality that is defined solely in terms of $(F,\delta)$ and applies equally to Euclidean, high-dimensional, and non-Euclidean data.
@@
Estimating and using $r_{F}(\cdot)$ in practice, however, raises two fundamental challenges. First, the induced pairwise preferences need not be transitive: it is possible to have $x$ preferred to $y$, $y$ preferred to $z$, and $z$ preferred to $x$, even at the population level. Second, empirical estimates of $p_{F}(x,y)$ are noisy and strongly dependent, as they are constructed from shared reference points. As a result, naive aggregation schemes, such as empirical win rates, can yield unstable rankings and provide limited inferential structure.
@@
To address these issues, we propose to aggregate induced preferences through a calibrated one-dimensional projection. Specifically, we consider a Bradley–Terry–Luce (BTL) type representation (Bradley and Terry, 1952; Luce, 1959) that maps pairwise preferences onto a transitive score scale by minimizing a population cross-entropy risk. Importantly, this formulation is used as a projection rather than as a correctly specified model: the resulting score function is defined as the unique minimizer of a well-posed M-estimation problem and exists regardless of whether the BTL form exactly matches the underlying preference probabilities. This projection enforces global coherence of the ordering, regularizes noisy comparisons, and yields interpretable log-odds scores that support efficient computation and principled out-of-sample evaluation.
@@
We call the resulting framework CORE (Centrality Ordering via Relative Evaluation). At the population level, we show that the calibrated score is a strictly monotone transformation of $r_{F}(\cdot)$, ensuring that the projection preserves the intrinsic preference-based ordering while providing a stable and interpretable coordinate system. At the sample level, we develop two complementary estimators: a convex optimization-based estimator and a fast spectral estimator based on the stationary distribution of an empirical comparison operator. We establish identifiability, consistency, and stability of the proposed scores under mild conditions, and show that, in settings where such structure is available, the induced population score satisfies the canonical center–outward properties associated with statistical depth. Through simulations and real-data analyses, we demonstrate that CORE yields sharper discrimination and more stable rankings than direct win-rate aggregation and widely used depth-based methods, particularly in high-dimensional and non-Euclidean settings.
@@
The remainder of the paper is organized as follows. Section 2 introduces the CORE framework and the proposed estimators. Section 3 develops the theoretical properties of the population and sample-level scores. Section 4 presents simulation studies. Section 5 concludes with a discussion and directions for future work.
@@ 2.1 From induced preferences to a calibrated centrality score
We adopt a BTL type formulation as a projection device. Let $\theta:\mathcal{X}\to\mathbb{R}$ be a real-valued score function and define
@@
We formalize this projection through a population M-estimation problem. Specifically, we quantify the discrepancy between $p_{F}$ and $q_{\theta}$ using the cross-entropy risk
@@
Since $q_{\theta}(x,y)$ depends only on score differences $\theta(y)-\theta(x)$, the risk $\mathcal{L}_{F}(\theta)$ is invariant under additive shifts of $\theta$. To ensure identifiability, we work with centered score functions. Let $L^{2}_{0}(F)=\{\theta\in L^{2}(F):\mathbb{E}_{F}[\theta(X)]=0\}$ and let $\Theta(F)\subset L^{2}(F)$ be a user-specified function class. We define $\Theta_{0}(F)=\Theta(F)\cap L^{2}_{0}(F)$ and set
@@
The resulting $\theta^{\star}$ is a population-level calibrated centrality score. For convenience, we also define the associated strength function $s^{\star}(x)=\exp\{\theta^{\star}(x)\}$, which induces the same ordering as $\theta^{\star}$.
@@ 2.2 Sample-level M-estimation of centrality scores
Let $x_{1},\ldots,x_{n}$ be an i.i.d. sample from $F$. Rather than imposing additional structure on $\theta(\cdot)$, we adopt a fully nonparametric sample-level formulation by estimating the score function only at the observed data points. Specifically, we seek to estimate the score vector $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{n})^{\top}$, where $\theta_{i}$ represents $\theta(x_{i})$.
Under the BTL projection (2), the probability that $x_{j}$ is preferred to $x_{i}$ is given by $q_{\theta}(x_{i},x_{j})=\sigma(\theta_{j}-\theta_{i})$, where $\sigma(t)=(1+e^{-t})^{-1}$. Replacing population quantities in (3) with empirical counterparts yields the sample objective
$\mathcal{L}_{n}(\boldsymbol{\theta})=\sum_{1\leq i<j\leq n}\Big[-\hat{p}_{ij}\log\sigma(\theta_{j}-\theta_{i})-(1-\hat{p}_{ij})\log\{1-\sigma(\theta_{j}-\theta_{i})\}\Big],$ (5)
where $\hat{p}_{ij}$ denotes an empirical estimator of $p_{F}(x_{i},x_{j})$. As in the population case, the objective is invariant under global shifts of $\boldsymbol{\theta}$, and we impose the centering constraint $\mathbf{1}^{\top}\boldsymbol{\theta}=0$. The estimator is defined as
$\widehat{\boldsymbol{\theta}}=\operatorname*{argmin}_{\boldsymbol{\theta}\in\mathbb{R}^{n},\ \mathbf{1}^{\top}\boldsymbol{\theta}=0}\mathcal{L}_{n}(\boldsymbol{\theta}).$ (6)
We solve (6) using projected first-order optimization; Algorithm 1 summarizes the procedure. The estimated strength scores are $\widehat{s}_{i}=\exp(\widehat{\theta}_{i})$, and the sample preference center is defined as $\widehat{\mu}=x_{\hat{\imath}}$ with $\hat{\imath}\in\arg\max_{i}\widehat{\theta}_{i}$.
@@
For large $n$, directly solving (6) can be computationally demanding. As a scalable alternative, we consider a spectral approximation that is closely related to classical spectral ranking methods for paired comparisons (Negahban et al., 2017). Given the empirical preference matrix $\{\hat{p}_{ij}\}$, we construct a Markov transition matrix $\mathbf{T}=[T_{ij}]_{i,j=1}^{n}$ defined by
@@
This construction induces a random walk that, at each step, moves from state $i$ to state $j$ with probability proportional to the probability that $j$ is preferred to $i$. The stationary distribution $\widehat{\mathbf{s}}$ of $\mathbf{T}$ (i.e., the probability vector satisfying $\widehat{\mathbf{s}}=\mathbf{T}^{\top}\widehat{\mathbf{s}}$) provides a strength score that serves as a spectral approximation to the BTL M-projection, in the sense that it replaces the global convex optimization in (6) with a linear eigenvector computation while preserving the same underlying preference aggregation structure. Centering the logarithm of $\widehat{\mathbf{s}}$ yields an approximate solution to (6). Algorithm 2 details the implementation. For faster computation, we also implement a quasi-Newton solver based on L-BFGS-B updates; the corresponding procedure is provided in Supplementary Section S.2 (Algorithm S.1).
@@
Input: pairwise preference matrix $\mathbf{P}=\{\hat{p}_{ij}\}_{1\leq i,j\leq n}$, stepsize $\eta>0$, tolerance $\varepsilon>0$, maximum iteration $T_{\max}$.
@@
Input: pairwise preference matrix $\mathbf{P}=\{\hat{p}_{ij}\}_{i,j=1}^{n}$ with $\hat{p}_{ii}=0$, tolerance $\varepsilon>0$, maximum iterations $T_{\max}$.
Step I (Transition matrix). Construct $\mathbf{T}=[T_{ij}]_{i,j=1}^{n}$ by
@@ 2.3 Construction of empirical pairwise preferences
Both estimators require empirical pairwise preferences $\hat{p}_{ij}$. Our primary construction is the leave-two-out estimator
$\hat{p}_{ij}^{\mathrm{lto}}=\frac{1}{n-2}\sum_{\ell\neq i,j}\mathbb{I}\{\delta(x_{\ell},x_{i})>\delta(x_{\ell},x_{j})\},$ (7)
which uses all available observations as reference points. This estimator is statistically efficient but induces dependence across entries of the preference matrix.
For theoretical analysis and uncertainty quantification, we also consider a data-reference estimator based on an independent sample $Z_{1},\ldots,Z_{m}\overset{\mathrm{i.i.d.}}{\sim}F$:
@@
Finally, sample splitting can be used to ensure conditional independence of the summands at the cost of evaluating centrality only on a subset of points.
2.4 Out-of-sample scoring
To extend the fitted scores beyond the observed sample, we employ kernel smoothing on the dissimilarity space. Given a kernel $K:[0,\infty)\to[0,\infty)$ and bandwidth $h>0$, we define weights
@@ 2.4 Out-of-sample scoring
This construction can be viewed as a Nadaraya–Watson-type extension on $(\mathcal{X},\delta)$. In practice, we use the Gaussian kernel $K(t)=\exp(-t^{2})$ and set $h$ to the median pairwise distance among the sample points.
@@
At the population level, the centrality score $\theta^{\star}(\cdot)$ is defined as a function on $\mathcal{X}$ through the M-projection in (4). At the sample level, we adopt a fully nonparametric formulation by estimating this score only at the observed data points, yielding a finite-dimensional vector $\widehat{\boldsymbol{\theta}}=(\widehat{\theta}_{1},\ldots,\widehat{\theta}_{n})^{\top}$.
This discretized estimation strategy avoids imposing additional structural assumptions on $\theta(\cdot)$ and aligns with classical depth- and rank-based constructions, which define population functionals but operate on sample-level scores. Out-of-sample values are subsequently obtained through smoothing, as described in Section 2.4.
An alternative approach is to estimate $\theta(\cdot)$ directly within a parametric or reproducing-kernel class from reference-based comparisons. Recent trainable centrality and ranking frameworks pursue this direction; see, for example, Vu et al. (2025). We view such parametric approaches as complementary, offering amortized out-of-sample scoring at the cost of additional modeling assumptions.
@@
A direct sample analog of (1) is the average win rate
@@
which can be viewed as the average win probability of $x_{j}$ against a randomly chosen opponent. Theorem 2 shows that, at the population level, the M-projection score $\theta^{\star}(\cdot)$ is a strictly monotone transform of $r(\cdot)$. As a result, $\hat{r}_{j}$ and $\widehat{\theta}_{j}$ induce aligned one-dimensional orderings in large samples.
However, the BTL projection provides a calibrated score representation that enforces global coherence of the ordering, regularizes noisy pairwise comparisons, and yields interpretable log-odds differences, which are not available from raw win rates.

### v2

@@ Abstract
Ranking or assessing centrality in multivariate and non-Euclidean data is difficult because there is no canonical order and many depth notions become computationally fragile in high-dimensional or structured settings. We introduce a preference-based notion of centrality defined through population proximity comparisons with respect to a random reference draw, yielding a metric-intrinsic statistical functional that is well-defined on general metric spaces. Because the induced pairwise preferences may be non-transitive, we map them to a coherent one-dimensional score via a Bradley–Terry–Luce cross-entropy projection, viewed as a calibrated aggregation device rather than a correctly specified model. We develop two finite-sample estimators a convex M-estimator and a fast spectral estimator based on a comparison operator, and establish identifiability and consistency under mild conditions. Simulations and real-data examples, including high-dimensional and functional observations, illustrate that the proposed scores provide stable, interpretable rankings aligned with the underlying preference centrality.
Keywords: centrality; ranking; statistical depth; metric spaces; paired comparisons; Bradley–Terry–Luce projection; spectral ranking
@@ 1 Introduction
Centrality is a fundamental concept in data analysis. It underlies representative point selection, outlier detection, risk stratification, and the construction of interpretable orderings. In univariate settings, the real line provides a canonical order, so centrality and ranking are naturally defined through ranks, quantiles, and order statistics. In contrast, modern data are frequently multivariate or live on non-Euclidean spaces such as manifolds, networks, graphs, and learned embedding spaces, where no intrinsic notion of order is available. In these settings, defining population-level centrality and estimating a stable, interpretable ordering from finite samples remain challenging problems. In many contemporary applications, the only primitive is a dissimilarity or distance oracle (possibly task-specific or learned), and there may be no meaningful linear, convex, or coordinate structure to exploit.
A broad literature has sought to generalize ranking and centrality beyond the univariate case. Classical approaches include componentwise and spatial ranks (Bickel, 1965; Hallin and Puri, 1995; Puri and Sen, 1966; Chaudhuri, 1996; Oja, 2010), depth-induced orderings (Liu and Singh, 1993; Zuo and Serfling, 2000), Mahalanobis and metric-based constructions (Hallin and Paindaveine, 2002, 2006; Pan et al., 2018), and more recent transport- or graph-based formulations (Deb and Sen, 2023; Zhou and Chen, 2023). Statistical depth, in particular, has emerged as a principled framework for multivariate centrality, formalizing center–outward ordering through axioms such as affine equivariance, maximality, monotonicity, and robustness (Tukey, 1974; Liu, 1990; Vardi and Zhang, 2000; Zuo, 2003; Parzen, 1962). Depth notions have also been developed for functional data and come with accompanying theoretical guarantees (Yeon et al., 2025; Bočinec et al., 2025). Despite their success, many existing depth notions rely on linear or convex structure and can lose resolution or become computationally burdensome in high dimensions or in complex geometries, where distance concentration, hubness, and anisotropy obscure purely geometric notions of centrality (Beyer et al., 1999; Radovanović et al., 2010; Feldbauer and Flexer, 2019). More broadly, when only a general dissimilarity is available, it is often unclear what population object a given surrogate score targets, and how pairwise information across the sample should be coherently aggregated into a global ordering.
@@
Throughout, we focus on the no-ties regime where $p_{F}(x,y)+p_{F}(y,x)=1$ for all distinct $x,y\in\mathcal{X}$. Intuitively, $r_{F}(y)$ is the probability that $y$ “wins” against a random opponent when judged by a random reference point; it coincides with the population counterpart of an average win-rate aggregation in paired-comparison settings. This functional defines a population notion of centrality determined solely by $(F,\delta)$ and is directly meaningful in Euclidean, high-dimensional, and non-Euclidean settings.
@@
Estimating and using $r_{F}(\cdot)$ in practice, however, raises two fundamental challenges. First, the induced pairwise preferences need not be transitive: it is possible to have $x$ preferred to $y$, $y$ preferred to $z$, and $z$ preferred to $x$, even at the population level. Second, empirical estimates of $p_{F}(x,y)$ are noisy and strongly dependent, as they are constructed from shared reference points. As a result, naive aggregation schemes, such as empirical win rates, can yield unstable rankings and do not provide a calibrated score scale with a likelihood-based inferential handle.
@@
To address these issues, we aggregate induced preferences through a calibrated one-dimensional projection. Specifically, we consider a Bradley–Terry–Luce (BTL) type representation (Bradley and Terry, 1952; Luce, 1959) that maps pairwise preferences onto a transitive score scale by minimizing a population cross-entropy risk. Importantly, this formulation is used as a projection rather than as a correctly specified model: the resulting score function is defined as the unique minimizer of a well-posed M-estimation problem and exists regardless of whether the BTL form exactly matches the underlying preference probabilities. This projection enforces global coherence of the ordering, regularizes noisy comparisons, and yields interpretable log-odds differences that provide a coherent and computationally tractable score scale.
@@
We call the resulting framework CORE (Centrality Ordering via Relative Evaluation). At the population level, we show that the calibrated score is a strictly monotone transformation of $r_{F}(\cdot)$, ensuring that the projection preserves the intrinsic preference-based ordering while providing a stable coordinate system. At the sample level, we develop two complementary estimators: a convex optimization-based estimator and a fast spectral estimator based on the stationary distribution of an empirical comparison operator. We establish identifiability, consistency, and stability of the proposed scores under mild conditions. In Euclidean settings where a center–outward geometry is meaningful, we further show that, under standard regularity conditions, the induced population score satisfies the canonical axioms associated with statistical depth. Through simulations and real-data analyses, we illustrate that CORE produces stable rankings aligned with the underlying preference centrality across high-dimensional and non-Euclidean settings.
@@
The remainder of the paper is organized as follows. Section 2 introduces the CORE framework and the proposed estimators. Section 3 develops the theoretical properties of the population and sample-level scores. Section 4 presents simulation studies and real data analysis is performed in Section 5. Section 6 concludes with a discussion and directions for future work.
@@ 2.1 From induced preferences to a calibrated centrality score
We adopt a BTL-type formulation as a projection device. Let $\theta:\mathcal{X}\to\mathbb{R}$ be a real-valued score function and define
@@
We formalize this projection through a population M-estimation problem. Since $p_{F}(x,y)\in(0,1)$ represents a comparison probability, a natural way to project possibly non-transitive preferences onto a parametric comparison scale is via a strictly proper scoring rule. We therefore quantify the discrepancy between $p_{F}$ and $q_{\theta}$ using the cross-entropy risk
@@
Up to an additive term that does not depend on $\theta$, minimizing $\mathcal{L}_{F}(\theta)$ is equivalent to minimizing the expected Kullback–Leibler divergence from the Bernoulli distribution with success probability $p_{F}(X,Y)$ to its logistic approximation $q_{\theta}(X,Y)$, averaged over $(X,Y)\sim F$. This defines a well-posed projection of the preference kernel onto the logistic comparison family.
Since $q_{\theta}(x,y)$ depends only on score differences $\theta(y)-\theta(x)$, the risk $\mathcal{L}_{F}(\theta)$ is invariant under additive shifts of $\theta$. To ensure identifiability, we work with centered score functions. Let $L^{2}(F)$ denote the space of measurable functions $\theta:\mathcal{X}\to\mathbb{R}$ with $\int\theta(x)^{2}\,dF(x)<\infty$, and let $L^{\infty}(F)$ denote the space of essentially bounded measurable functions. Define
$L^{2}_{0}(F)=\{\theta\in L^{2}(F):\mathbb{E}_{F}[\theta(X)]=0\}.$
Let $\Theta(F)\subset L^{2}(F)$ be a user-specified function class. In the most general case one may take $\Theta(F)=L^{2}(F)\cap L^{\infty}(F)$ to ensure square integrability and boundedness; more structured classes (e.g., RKHS or parametric families) can also be considered. We define $\Theta_{0}(F)=\Theta(F)\cap L^{2}_{0}(F)$ and set
@@
The resulting $\theta^{\star}$ is a population-level calibrated centrality score. As shown in Section 3, $\theta^{\star}(\cdot)$ preserves the intrinsic ranking induced by $r_{F}(\cdot)$ up to a strictly monotone transformation. For convenience, we also define the associated strength function $s^{\star}(x)=\exp\{\theta^{\star}(x)\}$, which induces the same ordering as $\theta^{\star}$.
@@ 2.2 Sample-level M-estimation of centrality scores
Let $x_{1},\ldots,x_{n}$ be an i.i.d. sample from $F$. Recall that the population target is the function $\theta^{\star}(\cdot)$ defined through the M-projection in (4). In practice, centrality is used to rank the observed sample points. Accordingly, rather than estimating $\theta^{\star}(\cdot)$ over the entire space $\mathcal{X}$, we approximate it by estimating its values at the observed data points, leading to a finite-dimensional score vector $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{n})^{\top}$, where $\theta_{i}$ is intended to estimate $\theta^{\star}(x_{i})$.
To construct an estimator consistent with the population projection, we replace the population risk $\mathcal{L}_{F}(\theta)$ by its empirical counterpart. Under the BTL representation (2), the comparison probability between $x_{i}$ and $x_{j}$ is $q_{\theta}(x_{i},x_{j})=\sigma(\theta_{j}-\theta_{i})$, where $\sigma(t)=(1+e^{-t})^{-1}$. Let $\widehat{\mathbf{P}}=\{\hat{p}_{ij}\}_{i,j=1}^{n}$ denote an empirical estimate of the preference kernel $p_{F}(x_{i},x_{j})$. The sample objective is then defined as
$\mathcal{L}_{n}(\boldsymbol{\theta};\widehat{\mathbf{P}})=\sum_{1\leq i<j\leq n}\Big[-\hat{p}_{ij}\log\sigma(\theta_{j}-\theta_{i})-(1-\hat{p}_{ij})\log\{1-\sigma(\theta_{j}-\theta_{i})\}\Big],$ (5)
which is the empirical analogue of (3).
As in the population case, the empirical objective depends only on score differences and is therefore invariant under global shifts of $\boldsymbol{\theta}$. We impose the centering constraint $\mathbf{1}^{\top}\boldsymbol{\theta}=0$ to ensure identifiability and define the M-estimator
$\widehat{\boldsymbol{\theta}}=\operatorname*{argmin}_{\boldsymbol{\theta}\in\Theta_{n}}\mathcal{L}_{n}(\boldsymbol{\theta};\widehat{\mathbf{P}}),\qquad\Theta_{n}:=\{\boldsymbol{\theta}\in\mathbb{R}^{n}:\mathbf{1}^{\top}\boldsymbol{\theta}=0\}.$ (6)
The objective is convex in $\boldsymbol{\theta}$, and strictly convex on $\Theta_{n}$ under mild connectivity conditions on the comparison graph induced by $\widehat{\mathbf{P}}$, ensuring uniqueness of the solution. We compute $\widehat{\boldsymbol{\theta}}$ using projected first-order optimization; Algorithm 1 summarizes the procedure. For faster optimization of (6), we also implement a quasi-Newton solver based on L-BFGS-B updates (Supplementary Section S.2).
The estimated strength scores are $\widehat{s}_{i}=\exp(\widehat{\theta}_{i})$, which approximate the population strengths $s^{\star}(x_{i})=\exp\{\theta^{\star}(x_{i})\}$. We define the sample preference center as
$\widehat{\mu}=x_{\hat{\imath}},\qquad\hat{\imath}\in\arg\max_{1\leq i\leq n}\widehat{\theta}_{i}.$
@@
For large $n$, solving (6) can be computationally demanding. As a scalable alternative, we construct a spectral approximation based on a comparison-induced Markov chain, closely related to classical spectral ranking methods for paired comparisons (Negahban et al., 2017).
Given the empirical preference matrix $\widehat{\mathbf{P}}=\{\hat{p}_{ij}\}_{i,j=1}^{n}$, define the transition matrix $\mathbf{T}=[T_{ij}]_{i,j=1}^{n}$ by
@@
This construction defines a random walk on the sample points in which transitions favor preferred alternatives. Under mild irreducibility conditions on $\widehat{\mathbf{P}}$, $\mathbf{T}$ admits a unique stationary distribution $\widehat{\mathbf{s}}$ satisfying $\widehat{\mathbf{s}}=\mathbf{T}^{\top}\widehat{\mathbf{s}}$.
We use $\widehat{\mathbf{s}}$ as a spectral strength score and define centered log-scores
$\widehat{\theta}_{i}=\log\widehat{s}_{i}-\frac{1}{n}\sum_{k=1}^{n}\log\widehat{s}_{k},\qquad i=1,\ldots,n,$
where $\widehat{s}_{i}$ denotes the $i$th entry of $\widehat{\mathbf{s}}$. This procedure replaces the global convex optimization in (6) by an eigenvector computation while preserving the same preference aggregation structure. Algorithm 2 details the implementation.
@@
Input: pairwise preference matrix $\widehat{\mathbf{P}}=\{\hat{p}_{ij}\}_{i,j=1}^{n}$, stepsize $\eta>0$, tolerance $\varepsilon>0$, maximum iteration $T_{\max}$.
@@
Input: pairwise preference matrix $\widehat{\mathbf{P}}=\{\hat{p}_{ij}\}_{i,j=1}^{n}$ with $\hat{p}_{ii}=0$, tolerance $\varepsilon>0$, maximum iterations $T_{\max}$.
Step I (Transition matrix). Construct $\mathbf{T}=\{T_{ij}\}_{i,j=1}^{n}$ by
@@ 2.3 Construction of empirical pairwise preferences
Both estimators require empirical approximations of the population preference probabilities $p_{F}(x_{i},x_{j})$.
Our primary construction is the leave-two-out estimator
$\hat{p}_{ij}^{\mathrm{lto}}=\frac{1}{n-2}\sum_{\ell\neq i,j}\mathbb{I}\{\delta(x_{\ell},x_{i})>\delta(x_{\ell},x_{j})\},\qquad i\neq j.$ (7)
Under the no-ties assumption and the i.i.d. sampling scheme,
$\mathbb{E}\!\left[\hat{p}_{ij}^{\mathrm{lto}}\mid x_{i},x_{j}\right]=p_{F}(x_{i},x_{j}),$
so the estimator is conditionally unbiased. It uses all remaining observations as reference points and is statistically efficient in finite samples. However, because different pairs share common reference observations, the entries of the empirical preference matrix $\widehat{\mathbf{P}}$ are dependent.
We also consider a data-reference estimator based on an independent sample $Z_{1},\ldots,Z_{m}\overset{\mathrm{i.i.d.}}{\sim}F$:
@@
Conditionally on $x_{i}$ and $x_{j}$, the indicators $\mathbb{I}\{\delta(Z_{l},x_{i})>\delta(Z_{l},x_{j})\}$, $l=1,\ldots,m$, are i.i.d. Bernoulli random variables with success probability $p_{F}(x_{i},x_{j})=\mathbb{P}_{Z\sim F}\{\delta(Z,x_{i})>\delta(Z,x_{j})\}$. Consequently,
$\mathbb{E}\!\left[\hat{p}_{ij}^{\mathrm{ref}}\mid x_{i},x_{j}\right]=p_{F}(x_{i},x_{j}),\qquad\mathrm{Var}\!\left(\hat{p}_{ij}^{\mathrm{ref}}\mid x_{i},x_{j}\right)=\frac{p_{F}(x_{i},x_{j})\{1-p_{F}(x_{i},x_{j})\}}{m}.$
Different pairs remain dependent because they share the same reference sample $\{Z_{l}\}_{l=1}^{m}$, but this construction separates the reference sample from the comparison sample and avoids the leave-two-out dependence induced by reusing $\{x_{\ell}\}$ as reference points.
Alternatively, sample splitting may be used to form disjoint comparison and reference sets from the original data. While this reduces the effective sample size, it can be convenient in practice.
Remark 1 (Optional out-of-sample extension).
The CORE scores are estimated on the observed sample points $\{x_{i}\}_{i=1}^{n}$, which is sufficient for ranking the sample. If one wishes to evaluate a score at a new point $x\in\mathcal{X}$, a simple extension is to smooth the fitted scores on the dissimilarity space via kernel weighting. Given a kernel $K:[0,\infty)\to[0,\infty)$ and bandwidth $h>0$, define
@@
This yields a Nadaraya–Watson-type extension on $(\mathcal{X},\delta)$. Common choices include a Gaussian kernel $K(t)=\exp(-t^{2})$ with a bandwidth selected from the empirical distribution of pairwise distances (e.g., the median), though other choices are possible.
@@
At the population level, $\theta^{\star}(\cdot)$ is defined on $\mathcal{X}$ via the M-projection (4). At the sample level, we estimate only its evaluations at the observed data points, yielding $\widehat{\boldsymbol{\theta}}\in\mathbb{R}^{n}$. This discretized strategy avoids imposing additional structural assumptions on $\theta(\cdot)$ and parallels classical depth and rank constructions, which define population functionals but operate on finite samples.
An alternative approach is to estimate $\theta(\cdot)$ directly within a parametric or reproducing-kernel class from pairwise comparisons (Vu et al., 2025). Such methods offer amortized out-of-sample scoring but require stronger modeling assumptions. We view these approaches as complementary to the nonparametric M-projection framework developed here.
@@
A direct sample analogue of (1) is the average win rate
@@
which represents the empirical probability that $x_{j}$ defeats a randomly chosen opponent. At the population level, Theorem 2 shows that the M-projection score $\theta^{\star}(\cdot)$ is a strictly monotone transformation of $r_{F}(\cdot)$. Consequently, $\hat{r}_{j}$ and $\widehat{\theta}_{j}$ induce asymptotically aligned orderings.
The M-projection, however, yields a globally coherent score representation that regularizes noisy pairwise comparisons and provides interpretable log-odds differences, features not available from raw win rates.